How do you change the distance between the plates of a capacitor?

thank you for your help!
Answer #1

We can change distance between two parallel plates of capacitor by maintaining same Potential Difference across the plates.

This can be done by changing the emf of power supply sources.

C= Q1/V2= Q2/V2 = Constant.

In this manner we can change distance between plats of parallel plates without effing their capcitance.